Cullumber Monograms sells stadium blankets that have been monogrammed with high school and university emblems. The blankets retail for $42 throughout the country to loyal alumni of over 3,300 schools. Cullumber’s variable costs are 42% of sales; fixed costs are $116,000 per month.

Assume that variable costs increase to 45% of the current sales price and fixed costs increase by $14,300 per month. If Cullumber were to raise its sales price by 11% to cover these new costs, what would be the new annual breakeven point in sales dollars? (Round sales price to 2 decimal places, e.g. 52.75 and final answer to 0 decimal places, e.g. 5,275.)

Breakeven sales
Answer #1

New variable cost = $42*45% = $18.9

New fixed cost = $116000 + $14300 = $130300

Fixed cost annual = $130300*12 = $1563600

New selling price = $42 + 11% = $46.62

New contribution = $46.62 - $18.90 = $27.72

New breakeven units = $1563600/$27.72 = 56407 units

New break even sales in $ = $56407 * 46.62 = $2629694

Feel free to ask any queries..

Also plz upvote it means a lot . Thank you